# General Exploratory Bonus for Optimistic Exploration in RLHF

### v1

@@ 4 General Exploratory Bonus with Optimism Principle
Motivated by the failure of the existing exploratory bonus, we now propose a novel framework, General Exploratory Bonus (GEB), and prove that it achieves optimism. We further show that prior exploratory bonuses—and broader variants—emerge as special cases of our formulation.
@@
As shown in Section 3, the failure of existing bonuses arises because the divergence constraints in $\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ force $\pi$ to align with $\pi_{\textrm{ref}}$, biasing exploration toward high-$\pi_{\text{ref}}$ regions. To achieve optimistic exploration, the optimal $\pi$ must instead counteract this regularization and move away from $\pi_{\textrm{ref}}$. Our key idea is to introduce an additional reference-dependent regulation into the reward, which offsets the influence of divergence regularization. The resulting exploratory bonus takes the form $\kappa\max_{\pi}J_{\beta,f}(\pi,R(r,\pi_{\textrm{ref}}))$. Note that the formulation of $R(\cdot,\cdot)$ can be diverse. Since the reward $R(r,\pi_{\textrm{ref}})$ is now explicitly dependent on $\pi_{\textrm{ref}}$, the inner optimal policy of the bi-level problem is $\frac{1}{Z_{R}(x)}\pi_{\textrm{ref}}(f^{\prime})^{-1}(\frac{R(r,\pi_{\textrm{ref}})}{\beta})$, where $Z_{R}(x)$ is a normalization term. Unlike previous cases, the optimal policy is no longer guaranteed to be positively correlated with $\pi_{\textrm{ref}}$, enabling the policy to deviate from the reference distribution.
As in Lemma 3.2, we can substitute the inner $\pi$ by the closed-form solution, and then utilize the reward reparameterization of $\alpha$-divergence (Wang et al., 2024) as $r=f^{\prime}(\pi/\pi_{\textrm{ref}})$ for a divergence instance $f$ to obtain the reparameterized exploratory bonus as
$\mathcal{L}_{\text{bonus}}(r(\pi))=\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(\cdot|x)}\Big[\frac{u}{Z_{R}(x)}f^{\prime}(u)-f(\frac{u}{Z_{R}(x)})\Big],$ (10)
where $u=(f^{\prime})^{-1}(R(\pi_{\textrm{ref}}(y|x),\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})/\beta)$ and $Z_{R}(x)=\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}u$. Since the domain of divergence class is generally $(0,+\infty)$ while there are no additional constraints on the formulation of $R$, $u$ can be flexibly formulated with $\pi$ and $\pi_{\textrm{ref}}$ unless $u\geq 0$.
@@
In our proposed exploratory bonus, the normalization term $Z_{R}(x)$ in Eq. 10 cannot be canceled. Fortunately, we prove the following lemma in Appendix E to show the equivalence between the two training objectives, one with and the other without $Z_{R}(x)$, which helps transform the objective to a succinct formulation for analyses and practical use.
@@
Denote two objectives as $h(u)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}uf^{\prime}(u)-f(u)$ and $\hat{h}(u)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{u}{Z(x)}f^{\prime}(u)-f(\frac{u}{Z(x)})$ where $u$ is a function with $\pi$ and $\pi_{\textrm{ref}}$. When $f^{\prime}(u)+uf^{\prime\prime}(u)-f^{\prime}(\frac{u}{Z(x)})]/[Z(x)uf^{\prime\prime}(u)]=\Lambda(x)$ is constant in $y$ and $\Lambda(x)>0$, minimizing the two objectives $\min_{\pi}-h(u)$ and $\min_{\pi}-\hat{h}(u)$ induce the same class of policies.
@@
Building on Lemma 4.1, we now prove that our proposed framework indeed achieves the optimism requirement.
@@
For each iteration of online RLHF, if the policy is updated by $\pi=\arg\max_{\pi}J_{\beta,f}(\pi,r)$ while its reward is trained with $r=\arg\min_{r}[\mathcal{L}_{BT}(\mathcal{D},r)-\kappa\mathcal{L}_{\text{bonus}}]$. When $f$ belongs to $\alpha$-divergence class, and $\mathcal{L}_{\text{bonus}}=\max_{\pi}J_{\beta,f}(\pi,R(r,\pi_{\textrm{ref}}))$, denote $u(\pi,\pi_{\textrm{ref}})=(f^{\prime})^{-1}\Big(R\Big((f^{\prime})^{-1}(\frac{\pi}{\pi_{\textrm{ref}}}),\pi_{\textrm{ref}}\Big)/\beta\Big)$, Then, $\frac{\partial^{2}\mathcal{L}_{\text{bonus}}}{\partial\pi\partial\pi_{\textrm{ref}}}\leq 0$ if $\forall(x,y);\frac{\partial u}{\partial\pi}+\pi_{\textrm{ref}}\frac{\partial^{2}u}{\partial\pi\partial\pi_{\textrm{ref}}}+\frac{(\alpha-1)\pi_{\textrm{ref}}}{u}\frac{\partial u}{\partial\pi}\frac{\partial u}{\partial\pi_{\textrm{ref}}}<0$ and $u>\alpha$, where $\alpha$ is a hyperparameter that defines $\alpha$-divergence (Definition 3.2).
$\mathcal{L}_{\text{bonus}}=\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\Big[uf^{\prime}(u)-f(u)\Big].$ (11)
For $\alpha$-divergences, the conditions in Lemma 4.1 are satisfied. Since $f^{\prime}(u)+uf^{\prime\prime}(u)-f^{\prime}(\frac{u}{Z})=u^{\alpha-1}(Z^{1-\alpha}-\alpha)/(1-\alpha)$ and $uf^{\prime\prime}(u)=u^{\alpha-1}$, the fraction $\Lambda(x)=(Z^{1-\alpha}-\alpha)/Z(1-\alpha)$ in Lemma 4.1 is independent of $y$. Since $u>\alpha$, $Z_{R}=\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}u>0$, thus $\Lambda(x)>0$ is also satisfied. Finally, the mixed second-order derivative of Eq. 11 is computed as
which achieves the optimism defined in Definition 3.1. $\square$
Table 2: General exploratory bonus under different divergence classes and design of $u$. Note all $u>\alpha$ when $0<\pi<1$. The presented bonus is simplified by removing constant coefficients or biases.
@@
In our formulation, $u$ can be flexibly defined in terms of $\pi,\pi_{\textrm{ref}}$ as long as it satisfies the derivative condition in Theorem 4.2 and $u>\alpha$. This flexibility highlights the extensibility of our framework. In particular, when $u$ depends only on $\pi$, any function with $u>\alpha$ and negative correlation with $\pi$ qualifies. In Table 2, we list several such choices of $u$, along with their corresponding reparameterized exploratory bonus under three different $\alpha$-divergences.
From a practical standpoint, since $\mathcal{L}_{\text{bonus}}$ is expressed as an expectation over $\pi_{\textrm{ref}}$, it does not require additional sampling and can be seamlessly integrated into iterative online RLHF. To avoid unintended decreases in the likelihood of preferred responses, however, we follow Chen et al. (2025) and restrict computation of the bonus on rejected responses to ensure that the probability of preferred responses continues to increase.
@@
Although we have shown that existing theoretical formulations of $\mathcal{L}_{\text{bonus}}$ fail to guarantee optimism, many practical implementations have nevertheless been effective through various approximations and adaptations. These approximations and adaptations are generally inextensible beyond the reverse KL divergence (detailed in Appendix B.2). In this subsection, we show that these practical implementations can be naturally subsumed into our GEB framework, and even broader objectives can be reinterpreted as instances of optimistic exploration. For example, Zhang et al. (2024a) and Xie et al. (2024) finally implement their exploratory bonus as $\kappa\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(y|x)}\log\pi(y|x)$, which belongs to GEB when $u(\pi)=-\log\pi+1$ and $f$ is KL-divergence. Similarly, Cen et al. (2025) implement the exploratory bonus as $\kappa\mathbb{E}_{x\sim\rho,y\sim\pi_{\text{cal}}(\cdot|x)}\log\frac{\pi}{\pi_{\textrm{ref}}}$ where $\pi_{\text{cal}}$ is a fixed calibration distribution. This also falls under GEB by setting $u=-\frac{\pi_{\text{cal}}}{\pi_{\textrm{ref}}}\log\frac{\pi}{\pi_{\textrm{ref}}}-\frac{\pi_{\text{cal}}}{\pi_{\textrm{ref}}}\log{\pi_{\textrm{ref}}}+1$ and $f$ is KL-divergence. The corresponding reparameterized exploratory bonus reduces to $\mathbb{E}_{x\sim\rho,y\sim\pi_{cal}}-\log\frac{\pi}{\pi_{\textrm{ref}}}+C(x)$, where $C(x)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}[-\frac{\pi_{\text{cal}}}{\pi_{\textrm{ref}}}\log{\pi_{\textrm{ref}}}+1]$. Interestingly, even objectives not explicitly designed for exploration can be reinterpreted through our GEB framework. For instance, Chen et al. (2025) augment the DPO loss with an additional term $\kappa\mathbb{E}_{x,y\sim\pi_{ref}}\sigma(-\beta\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})$, which was originally introduced to control sample complexity. In our framework, this corresponds to optimistic exploration with $u=-\sigma(-\beta\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})+1$.
@@ 5 Experiments
5.1 Experimental Settings
Baselines.
5.2 Results & analyses
The experimental results are shown in Table 3. Across both backbones, GEB generally outperforms f-DPO and FEB. Under the KL-divergence, GEB displays better or at least on-par performance compared to prior exploratory-bonus methods. Notably, the win-rate increases over 1.82% and 0.94% under the KL-divergence, over 2.36% and 1.29% under the Hellinger Distance, compared with their f-DPO counterpart. GPT-4 evaluation on the Alpaca benchmark also shows consistent performance gains on out-of-domain alignment task. While GEB maintains on par, or usually better results in MATH, showing less performance degradation beyond alignment, known as alignment tax (Noukhovitch et al., 2023; Lin et al., 2024).
@@ 5.2 Results & analyses
The choice of $\kappa$.
Since the formulation of $u$ in Eq. 10 is flexible, the scale of the GEB term can differ substantially across designs, hence the absolute value of the bonus is less informative. Instead, we examine the relative ratio of the bonus term to the vanilla RL loss $|\kappa\mathcal{L}_{bonus}|/|\mathcal{L}_{RL}|$, which provides a more consistent basis for comparison and offers better practical guidance for tuning $\kappa$ across diverse settings. As shown in Fig. 3, performance remains stable when the ratio lies within a suitable range (1e-2 to 1e-6 in our case). However, if the ratio is too large, it impedes optimization of the RL objective and degrades performance; if too small, the exploration incentive in uncertain regions diminishes and performance reverts to the vanilla baseline.

### v4

@@ 4 General Exploratory Bonus with Optimism Principle
Motivated by the failure of the existing optimistic exploration works, we now propose a novel framework, General Exploratory Bonus (GEB), and prove that it achieves optimism. We further show that prior heuristic bonuses—and their broader variants—emerge as special cases of our formulation.
@@
As shown in the previous section, existing bonus schemes fail because the divergence constraints in $\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ force the optimal policy $\pi$ to remain close to $\pi_{\textrm{ref}}$, thereby biasing exploration toward regions where $\pi_{\textrm{ref}}$ is large. Achieving optimistic exploration requires the optimal $\pi$ to counteract this effect and deviate from $\pi_{\textrm{ref}}$. Our key idea is therefore to incorporate an additional $\pi_{\textrm{ref}}$-dependent term into the reward that offsets the influence of the divergence regularization. The resulting exploratory bonus takes the form
$\mathcal{L}_{\text{bonus}}=\max_{\pi}J_{\beta,f}(\pi,R),$ (8)
where our new reward formulation, $R$, now depends not only on the original reward model $r(x,y)$ but also on $\pi_{\textrm{ref}}(y|x)$. Now, the optimal policy of $\max_{\pi}J_{\beta,f}(\pi,R(x,y))$ can be obtained by replacing $r(x,y)$ with $R(x,y)$ in Eq 6. This yields $\pi^{*}(y|x)=\frac{1}{Z_{R}(x)}\pi_{\textrm{ref}}(f^{\prime})^{-1}(\frac{R(x,y)}{\beta})$, where $Z_{R}(x)$ is a normalization term.
Following Lemma 3.2, we substitute $\pi(y|x)$ in $\max_{\pi}J_{\beta,f}(\pi,R(x,y))$ with its optimal form $\pi^{*}(y|x)$ and then apply the reward reparameterization trick for $\alpha$-divergences (Wang et al., 2024), i.e., $r(x,y)=f^{\prime}(\pi^{*}(y|x)\big/\pi_{\textrm{ref}}(y|x))$, where $f$ specifies the divergence. Given this policy-reparameterized reward, we specify the exploratory bonus term in Eq 8 as follows:
$\mathcal{L}_{\text{bonus}}=\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(\cdot|x)}\Big[\frac{u(x,y)}{Z_{R}(x)}f^{\prime}(u(x,y))-f(\frac{u(x,y)}{Z_{R}(x)})\Big].$ (9)
In Eq. 9, we introduced $u(x,y)$ as an atomic function employed to construct the actual loss, and it is given by $u(x,y)=(f^{\prime})^{-1}(R(x,y)/\beta)$ in this setup. Note that, after reward reparameterization, $u(x,y)$ can be expressed in terms of $\pi(y|x)$ and $\pi_{\textrm{ref}}(y|x)$. Moreover, as the functional form of $R(x,y)$ is not restricted to a specific class, $u(x,y)$ can be instantiated in many ways using these two distributions, while it must satisfy $u(x,y)>0$ for all $x,y$ to ensure that the argument of $f^{\prime}(\cdot)$ lies within its domain. See Table 2 for the example entries we are considering in this work.
@@
In our proposed exploratory bonus, the normalization term $Z_{R}(x)$ in Eq. 9 cannot be eliminated. Fortunately, Lemma 4.1 (proved in Appendix B.4) shows that the training objectives with and without $Z_{R}(x)$ are equivalent. This equivalence allows us to convert the objective into a more concise form, facilitating both analysis and practical implementation.
@@
Denote two objectives as $h(u(x,y))=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}u(x,y)f^{\prime}(u(x,y))-f(u(x,y))$ and $\hat{h}(u(x,y))=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{u(x,y)}{Z_{R}(x)}f^{\prime}(u)-f(\frac{u(x,y)}{Z_{R}(x)})$ where $u(x,y)$ is a function with $\pi(y|x)$ and $\pi_{\textrm{ref}}(y|x)$. If the ratio
$\frac{f^{\prime}(u(x,y))+u(x,y)f^{\prime\prime}(u(x,y))-f^{\prime}(\frac{u(x,y)}{Z_{R}(x)})}{Z_{R}(x)u(x,y)f^{\prime\prime}(u(x,y))}=\Lambda(x)$ (10)
is independent of $y$ and $\Lambda(x)>0$, then minimizing the two objectives, $\min_{\pi}-h(u(x,y))$ and $\min_{\pi}-\hat{h}(u(x,y))$, yields the same class of optimal policies.
Note that the $\alpha$-divergence ($0\leq\alpha\leq 1$; see Def. 3.2) naturally satisfies the condition in Eq. 10 whenever $u(x,y)>\alpha$. As we show in the next paragraph, enforcing $u(x,y)>\alpha$ is straightforward in practice, which grants our framework substantial flexibility and extensibility. Leveraging Lemma 4.1, we can thus rewrite our objective in Eq. 9 into a concise and analytically convenient form without the normalization term:
$\mathcal{L}_{\text{bonus}}=\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\Big[u(x,y)f^{\prime}(u(x,y))-f(u(x,y))\Big],$ (11)
where $u(x,y)$ is flexibly formulated by $\pi(y|x)$ and $\pi_{\textrm{ref}}(y|x)$ satisfying $u(x,y)>\alpha$.
@@
Building on Lemma 4.1, we now show that our proposed framework achieves the optimism condition in Definition 3.1 (See Appendix B.5 for the proof).
@@
Consider an $\alpha$-divergence $f$ with $0\leq\alpha\leq 1$, and the exploratory bonus $\mathcal{L}_{\text{bonus}}=\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\Big[u(x,y)f^{\prime}(u(x,y))-f(u(x,y))\Big]$, where $u(x,y)$ is a function dependent on $\pi(y|x)$ and $\pi_{\textrm{ref}}(y|x)$. For any $(x,y)$, if $\frac{\partial u}{\partial\pi}+\pi_{\textrm{ref}}\frac{\partial^{2}u}{\partial\pi\partial\pi_{\textrm{ref}}}+\frac{(\alpha-1)\pi_{\textrm{ref}}}{u}\frac{\partial u}{\partial\pi}\frac{\partial u}{\partial\pi_{\textrm{ref}}}<0$ and $u(x,y)>\alpha$, the optimism condition in Definition 3.1 is satisfied; that is, $\frac{\partial^{2}\mathcal{L}_{\text{bonus}}}{\partial\pi\partial\pi_{\textrm{ref}}}\leq 0$.
Table 2: GEB under different divergence classes and design of $u$. Note that (1) now the bonus term can be computed without the reference probability mass $\pi_{\textrm{ref}}$; (2) all of these $u$ instantiations meet the condition $u>\alpha$ when $0<\pi<1$. The presented bonuses are simplified by removing constants.
@@
In our formulation, $u(x,y)$ can be flexibly defined in terms of $\pi(y|x)$ and $\pi_{\textrm{ref}}(y|x)$ as long as it satisfies the derivative condition in Theorem 4.2 and $u(x,y)>\alpha$. In particular, when $u(x,y)$ depends solely on $\pi(y|x)$ and is independent of $\pi_{\textrm{ref}}(y|x)$, any function that is strictly decreasing in $\pi$ with $u(x,y)>\alpha$ constitutes a valid choice. This design flexibility underscores the extensibility of our framework. In Table 2, we list several such choices of $u$, along with their corresponding reparameterized exploratory bonus terms under three different $\alpha$-divergences. From a practical standpoint, since $\mathcal{L}_{\text{bonus}}$ is computed as an expectation over $\pi_{\textrm{ref}}(\cdot|x)$, it does not require additional sampling and can be seamlessly integrated into iterative online RLHF. Meanwhile, to avoid unintended decreases in the likelihood of preferred responses, we follow Chen et al. (2025) and restrict the computation of the bonus on rejected responses to ensure that the probability of preferred responses continues to increase.
@@
Although we have shown that existing theoretical formulations of $\mathcal{L}_{\text{bonus}}$ fail to guarantee optimism, many practical implementations have nevertheless been effective through various approximations and adaptations. These approximations and adaptations are generally inextensible beyond the reverse KL divergence (detailed in Appendix B.3). In this paragraph, we show that these practical implementations can be naturally subsumed into our GEB framework, and even broader objectives can be reinterpreted as instances of optimistic exploration. For example, Zhang et al. (2024a) and Xie et al. (2024) finally implement their exploratory bonus as $\kappa\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(y|x)}\log\pi(y|x)$, which belongs to GEB when $u=-\log\pi+1$ and $f$ is KL-divergence. Similarly, Cen et al. (2025) implement the exploratory bonus as $\kappa\mathbb{E}_{x\sim\rho,y\sim\pi_{\text{cal}}(\cdot|x)}\log\frac{\pi}{\pi_{\textrm{ref}}}$ where $\pi_{\text{cal}}$ is a fixed calibration distribution. This also falls under GEB by setting $u=-\frac{\pi_{\text{cal}}}{\pi_{\textrm{ref}}}\log\frac{\pi}{\pi_{\textrm{ref}}}-\frac{\pi_{\text{cal}}}{\pi_{\textrm{ref}}}\log{\pi_{\textrm{ref}}}+1$ and $f$ is KL-divergence. Interestingly, even objectives not explicitly designed for exploration can be reinterpreted through our GEB framework. For instance, Chen et al. (2025) augment the DPO loss with an additional term $\kappa\mathbb{E}_{x,y\sim\pi_{ref}}\sigma(-\beta\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})$, which was originally introduced to control sample complexity. In our framework, this corresponds to optimistic exploration with $u=-\sigma(-\beta\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})+1$.
Takeaways.
In summary, we introduce a general formulation of the exploratory bonus term, GEB, which showcases the following unique strengths: (1) In contrast to the prior bonus terms, GEB meets the optimism condition theoretically, thus provably boosts exploration on relatively untapped region; (2) GEB spans a broad class of instantiations as shown in Table 2, specified by the combination of the $\alpha$-divergence class and the functional form of $u$, which provides a plenty options for practitioners to choose on demand; (3) GEB is practical as it does not incur additional sampling costs and can be seamlessly integrated into the existing iterative online RLHF framework (Appendix C.1); and (4) GEB offers an unified understanding of existing methods by generalizing them as special cases in a single flexible formulation.
@@ 5 Experiments
5.1 Experimental Settings
Baselines.
5.2 Results & analyses
The experimental results are shown in Table 3. Across both backbones, GEB generally outperforms f-DPO and FEB. Under the KL-divergence, GEB displays better or at least on-par performance compared to prior exploratory-bonus methods. Notably, the win-rate increases over 1.82% and 0.94% under the KL-divergence, over 2.36% and 1.29% under the Hellinger Distance, compared with their f-DPO counterpart. GPT-4 evaluation on the Alpaca benchmark also shows consistent performance gains on the out-of-domain alignment task. While GEB maintains on par, or usually better results in MATH, showing less performance degradation beyond alignment, known as alignment tax (Noukhovitch et al., 2023; Lin et al., 2024).
@@ 5.2 Results & analyses
The choice of $\kappa$.
Since the formulation of $u$ in Eq. 9 is flexible, the scale of the GEB term can differ substantially across designs, hence the absolute value of the bonus is less informative. Instead, we examine the relative ratio of the bonus term to the vanilla RL loss $|\kappa\mathcal{L}_{bonus}|/|\mathcal{L}_{RL}|$, which provides a more consistent basis for comparison and offers better practical guidance for tuning $\kappa$ across diverse settings. As shown in Fig. 3, performance remains stable when the ratio lies within a suitable range (1e-2 to 1e-6 in our case). However, if the ratio is too large, it impedes optimization of the RL objective and degrades performance; if too small, the exploration incentive in uncertain regions diminishes, and performance reverts to the vanilla baseline.
@@
Table 6: The performance (Pass@16) of DPO and three GEB variants on math reasoning tasks.
MATH500 OlympiadBench AIME 2025 Avg.
DPO 89.80 57.78 23.23 56.94
GEB-$\pi$ 92.80 64.59 28.23 61.87
GEB-$1/\pi$ 93.00 65.78 29.48 62.75
GEB-$\text{arctanh}\pi$ 92.80 64.59 29.38 62.26
GEB improves exploration in challenging problem instances.
To further illustrate when GEB offers the greatest benefit, we conduct an additional experiment targeting prompts whose preferred (i.e., correct) answers lie in regions rarely explored by the reference model during sampling. Starting from the DAPO-MATH-17K dataset (Yu et al., 2025), we filter out prompts for which the reference model, Qwen2.5-7B (Qwen et al., 2025), can generate a correct answer within two samples at temperature 0.6. This removes 5K prompts whose correct answers fall in frequently visited regions, leaving 12K prompts whose correct answers are less likely to be sampled. We use an open-source verifier, rather than reward models, to determine correctness. Each preference pair consists of one correct response and one wrong response. We set the learning rate to 5e-7 with a warm-up ratio of 0.1. For the reported runs under KL divergence, the $\kappa$ values for the three GEB variants are 50, 5e-2, and 5e-6, respectively.
As shown in Table 6, all three GEB variants consistently surpass the DPO baseline with a clear and stable performance margin. Notably, the performance gain of GEB is more than 6% on OlympiadBench and AIME 2025. It indicates that our GEB can significantly enhance the exploration under scenarios where the correct answers lie in a region with a small $\pi_{\textrm{ref}}$, i.e., an underexplored region. An additional experiment is provided in Appendix D.